OpenStudy (anonymous):

A fence is to be built to enclose a rectangular area of 320 square feet. The fence along three sides is to be made of material that costs 3 dollars per foot, and the material for the fourth side costs 12 dollars per foot. Find the length L and width W (with W≤L) of the enclosure that is most economical to construct.

OpenStudy (anonymous):

if let L=W then L^2=320 => L=17.89 ft 2(L+L)=4L =(4*17.89) ft the cost of three sides= 3*(3*17.89)=161 dollars the fourth one =12*17.89=214.68 dollars